# A Relative Homology Theory of Representation in Neural Networks

### v2

@@ Abstract
Previous research has proven that the set of maps implemented by neural networks with a ReLU activation function is identical to the set of piecewise linear continuous maps. Furthermore, such networks induce a hyperplane arrangement splitting the input domain into convex polyhedra $G_{J}$ over which the network $\Phi$ operates in an affine manner.
In this work, we leverage these properties to define the equivalence class of inputs $\sim_{\Phi}$, which can be split into two sets related to the local rank of $\Phi_{J}$ and the intersections $\cap\text{Im}\Phi_{J_{i}}$. We refer to the latter as the overlap decomposition $\mathcal{O}_{\Phi}$ and prove that if the intersections between each polyhedron and the input manifold are convex, the homology groups of neural representations are isomorphic to relative homology groups $H_{k}(\Phi(\mathcal{M}))\simeq H_{k}(\mathcal{M},\mathcal{O}_{\Phi})$. This lets us compute Betti numbers without the choice of an external metric. We develop methods to numerically compute the overlap decomposition through linear programming and a union-find algorithm.
Using this framework, we perform several experiments on toy datasets showing that, compared to standard persistent homology, our relative homology-based computation of Betti numbers tracks purely topological rather than geometric features. Finally, we study the evolution of the overlap decomposition during training on various classification problems while varying network width and depth and discuss some shortcomings of our method.
Machine Learning, ICML
@@ 1 Introduction
Deep learning has proven to be an incredibly powerful framework for learning and generalizing on highly complex tasks. Despite its widespread success, our theoretical understanding of this framework is still heavily lacking. What we do know is that deep neural networks are universal approximators (Hornik, 1991) and therefore have unbounded expressive power. If we further restrict ourselves to neural networks with ReLU activation functions, we also know that they are exactly equal to the set of continuous piecewise-linear (CPWL) functions (Arora et al., 2016).
@@
A CPWL function applies a different affine function over different pieces of its domain. Naturally, this phenomenon is also observed in neural networks, where the structure of the weights splits the domain into different polyhedra called linear regions (Pascanu et al., 2013; Montufar et al., 2014). Quantifying the number (Arora et al., 2016; Raghu et al., 2017; Serra et al., 2018; Hanin & Rolnick, 2019) as well as other properties (Fan et al., 2023) of the linear regions that a network generates is thought to reflect the expressivity of a neural network and is a rich subfield in its own right.
@@
Another avenue of research that has shown a lot of promise in many fields of science is Topological Data Analysis (TDA), which attempts to quantify the topological features of data (Carlsson, 2009; Wasserman, 2018). In the context of deep learning, it has been used to relate the topology of training trajectories (Birdal et al., 2021; Dupuis et al., 2023) and network weights (Rieck et al., 2018; Gutiérrez-Fandiño et al., 2021; Andreeva et al., 2023) to the generalization error of a neural network. In addition, while counting linear regions is one measure of expressivity, another way to quantify it is to look at the topology of the decision boundary of a neural network (Guss & Salakhutdinov, 2018; Petri & Leitão, 2020; Grigsby & Lindsey, 2022) or the degree to which a network can change the topology of the input domain as it is propagated through its layers (Naitzat et al., 2020; Wheeler et al., 2021).
@@
While the examples above are only a small subset of the plethora of novel ideas for how to apply topological tools to study neural networks, many, if not most, such ideas are somehow based on persistent homology (Zomorodian & Carlsson, 2004). Despite the power of this method, it is known that it is sensitive to outliers, undersampling, and highly non-linear mappings. Furthermore, it is based on calculating distances between points and therefore requires the choice of a metric. This choice leads to the identification of not only topological features, but also geometric ones like curvature (Bubenik et al., 2020) and convexity (Turkes et al., 2022). This might be a benefit in some cases, but when reasoning solely about topological features, we would like to avoid contamination from geometric sources.
@@
These issues are especially relevant in the context of deep neural networks as they apply highly non-linear transformations of the input domain that increase curvature (Poole et al., 2016) in the output space, making their analysis incredibly difficult. To avoid this, we study topological features in the input space through a purely topological lens. We do this using relative homology (Hatcher, 2002), which except for a few works (Pokorny et al., 2016; Blaser & Brun, 2022; Beshkov & Einevoll, 2024), has received surprisingly little attention in the deep learning literature.
@@ 1.1 Our contributions
Instead of having to deal with an external space, induced by a map, in which the choice of a metric is ambiguous, we can imagine the following. Take a dataset, determine which pieces are glued to each other by the map, and then count the holes in the manifold after this gluing. Under some simple assumptions, these three steps can be rigorously identified with a manifold $\mathcal{M}$, a quotient space $\sim_{\Phi}$ and the relative homology groups $H_{k}(\mathcal{M},\sim_{\Phi})$ respectively (see Appendix A for more background on these terms). This procedure avoids the leakage of geometric information into our estimates of homology since it only depends on the quotient space $\sim_{\Phi}$. It does, however, require knowledge or at least a good estimate of the topology of the initial manifold $\mathcal{M}$ and a robust way to identify the quotient space $\sim_{\Phi}$.
In this work, we show how these steps can be carried out in the context of ReLU neural networks, further developing previous work relating polyhedral decompositions and homology (Liu et al., 2023). There are two difficulties with making such an approach rigorous. Firstly, how can we actually calculate the quotient space $\sim_{\Phi}$? Secondly, given this quotient space, how and when can we calculate the relative homology groups $H_{k}(\mathcal{M},\sim_{\Phi})$?
Our fundamental realization is that since ReLU neural networks split the input domain into convex polyhedra, we can separate the sources of gluing (also called non-injectivity) into the local rank of the map at each polyhedron and the non-trivial intersections between the images of different polyhedra. We refer to these as the rank and the overlap source, respectively, and show that the latter can be computed either approximately, using a distance threshold, or exactly, using feasibility checks through linear programming. Afterwards, we prove under which conditions the overlap decomposition is sufficient to calculate relative homology groups.
We then compare the performance of persistent and relative homology when it comes to determining the Betti numbers of toy datasets with a known topology. Our approach also gives a new perspective on how topological information propagates through the layers of a neural network. Using relative homology, we observe that topological simplification happens much more gradually than observed by previous persistent homology-based calculations (Naitzat et al., 2020). Finally, we study the properties of the overlap decomposition in randomly initialized and trained neural networks and show that the volume of overlapping regions decreases after training, whereas the number of overlap regions tends to increase.
@@ 2 Decompositions of Neural Networks
Before diving deeper into the mathematics, let us first fix our terminology. We will assume that a dataset $\mathcal{D}=\{x_{1},x_{2},...,x_{K}\}$ is sampled from some compact manifold $\mathcal{M}$, with or without boundary, embedded in $\mathbb{R}^{n_{0}}$. A neural network is a function $\Phi:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{1}}...\to\mathbb{R}^{n_{L}}$, where $n_{l}$ is the $l$-th layer of the network. Between each pair of layers, we have preactivations $T_{l}:\mathbb{R}^{n_{l-1}}\to\mathbb{R}^{n_{l}}$ given by the affine functions $T_{l}(z)=Wz+b$ and representations given by $S_{l}(z)=\text{ReLU}(T_{l}(z))$. We will denote the output at layer $l$ by $\Phi^{l}:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{l}}$ which is given by function composition $\Phi^{l}=S_{l}\circ S_{l-1}\circ...\circ S_{1}=\prod\limits_{k=1}\limits^{l}S_{k}$. We will use the convention that the last layer does not apply an activation function so $\Phi^{L}=T_{L}\circ\prod\limits_{k=1}\limits^{L-1}S_{k}$.
@@ 2.1 ReLU networks and linear regions
It is well known in the literature that every ReLU neural network is equivalent to a continuous piecewise-linear function and vice versa (Arora et al., 2016) and that such a network decomposes input space into convex polyhedra (Montufar et al., 2014; Balestriero et al., 2019; Zhang & Wu, 2020; Liu et al., 2023). This happens as a result of the fact that the affine functions $T_{l}$ can be interpreted as specifying conditions $CT_{l}$, where $C$ is some diagonal matrix with entries in $\{-1,1\}$ defining a polyhedron $P=\{x|CT_{l}(x)\leq 0\}$. Each subsequent layer further decomposes the input space into more polyhedra, since it adds more conditions to the previous decomposition.
@@
Then the local layer decomposition supported on $R$ is defined by the set going over all possible $J$’s and $\mathcal{L}^{l}|_{R}=\{L^{l}_{J}|_{R}\}$.
While the codewords within the same region of the local layer decomposition are constant, that does not mean that the network applies the same affine function over the region (see B.6 in the Appendix). However, this property does hold in a different decomposition called the polyhedral decomposition (Liu et al., 2023). Instead of looking at the codeword at layer $l$ we can stack together all codeword vectors from the previous layers $C_{l}(x)=[c_{1}(x),c_{2}(x),...,c_{l}(x)]$ forming a new global codeword vector. This leads us to the next decomposition,
@@
Then the polyhedral decomposition supported on $R$ is defined by the set going over all possible $J$’s and $\mathcal{G}^{l}|_{R}=\{G^{l}_{J}|_{R}\}$.
@@ 2.2 The overlap decomposition
The polyhedral decomposition is widely known in the literature, here we define a new type of decomposition that can be leveraged to compute the homology groups of a neural network, which we will call the overlap decomposition. Our approach is intuitively justified by the realization that if neural networks are equivalent to continuous piecewise-linear functions, then the only way that the topology of the input space can change is through non-injective transformations. Since the network operates piecewise, we can have non-injectivity from two sources; see Appendix B.2 for a proof:
@@
    The map $\Phi^{l}|_{G^{l}_{J}}$ is low-rank and projects the region $G^{l}_{J}$ to a lower-dimensional subspace. We will call this the rank source.
@@
    For a set of maps $\{\Phi^{l}|_{G^{l}_{J_{1}}},\Phi^{l}|_{G^{l}_{J_{2}}},...,\Phi^{l}|_{G^{l}_{J_{N}}}\}$ on different polyhedra, there is a non-trivial intersection of the images over their respective polyhedra $\bigcap\limits_{n}\text{Im}\Phi^{l}|_{G^{l}_{J_{n}}}\neq\emptyset$. We will call this the intersection source.
The first condition has been described in previous work (Beshkov & Einevoll, 2024), but as we will see later, it can be ignored when the intersections between the data manifold and the regions $G^{l}_{J}$ are convex. Therefore, we will focus on defining the overlap decomposition based on the second condition, with the agreement that one should be careful when this condition fails to hold. With this in mind, let us define the overlap decomposition.
@@
A neural network $\Phi^{l}$ has an the equivalence class which describes all non-injective regions of a network,
@@
The overlap decomposition describes the part of the equivalence class coming from the intersection source and is given by,
$\begin{split}&\mathcal{O}_{\Phi^{l}}=\{[x]|\Phi(y)\in\bigcap\limits_{i\in I}\Phi(G^{l}_{J_{i}})\text{ for }\\ &I\subset\{1,2,...,\text{\# polyhedra}\}\text{ and }\Phi(y)\not\in\bigcap\limits_{i\in K}\Phi(G^{l}_{J_{i}})\\ &\text{ for any }I\subsetneq K,\Phi(x)=\Phi(y),|I|>1\}.\end{split}$ (5)
This decomposition can be fully computed given that we have knowledge of the polyhedral decomposition $\mathcal{G}^{l}$, by identifying the intersections between all pairs of different polyhedra and using a union-find structure on top (Kleinberg, 2006). In the next section, we discuss how to approximate or exactly compute this decomposition for a neural network and a supporting dataset. We also discuss how it can be leveraged to compute the homology groups of a neural representation.
@@ 3 Relating the Overlap Decomposition and Homology
3.1 Numerical Determination of the Overlap Decomposition
We have stated that the overlap decomposition is a subset of the equivalence class $\sim_{\Phi}=\{[x]|\Phi(x)=\Phi(y)\}$. However, we have yet to show how to actually compute this set when working with finite data samples. Here we propose two methods of increasing precision which can be used to find the overlap decomposition. This comes at the cost of an increase in their algorithmic complexity.
3.1.1 Distance threshold
Even if parts of the underlying data manifold are glued together, when working with finite samples, it is highly unlikely that we will ever observe actual perfect equality between points. The distance threshold approach tries to avoid this issue by simply replacing the strict equivalence condition by $\leq_{\epsilon\Phi}=\{[x]|d(\Phi(x),\Phi(y))\leq\epsilon\}$. Pseudocode for this algorithm is shown in 1.
Algorithm 1 Distance threshold
  Input: The output of a model $\Phi_{l}(x)$ and a threshold $\epsilon$
  $D\leftarrow d(\Phi_{l}(x),\Phi_{l}(y))$ Compute distances
  $\mathcal{O}\leftarrow\arg[D\leq\epsilon]$ Identify close pairs
  $\mathcal{O}\leftarrow UF(\mathcal{O})$ Union-Find to get partially overlapping pairs
  Return $\mathcal{O}$
This method is still highly data-dependent and non-exact. Since it is evaluated in the output space it carries all of the problems that come with sparse data sampling and highly non-linear maps. Its advantage is that it is easy to understand and has a low computational cost.
3.1.2 Linear programming
As we will show later, under the assumption that the intersection between a polyhedron from the polyhedral decomposition and the data manifold $P\cap\mathcal{M}$ is a convex set (this is automatically true if $\mathcal{M}$ is convex), we can use the polyhedral decomposition to exactly determine the overlaps between data samples. For two populated polyhedra, meaning there exist points in the dataset $\mathcal{D}$ that live in each of these polyhedra, $P_{1},P_{2}\in\mathcal{G}^{l}$ the overlaps $\Phi_{l}(P_{1})\cap\Phi_{l}(P_{2})$ can be determined through linear programming. This is done by checking if given a point $p\in P_{1}$ there is another point $p^{\prime}\in P_{2}$ such that $\Phi_{l}(p)=\Phi_{l}(p^{\prime})$. This second point $p^{\prime}$ can be found through linear programming under the constraints given by the H-representation of $P_{1}$ and $P_{2}$. Thus, to check whether there is a point in $P_{2}$ that overlaps with $p\in P_{1}$, we get the following feasibility linear program,

### v3

# A Quotient Homology Theory of Representation in Neural Networks

@@ Abstract
Previous research has proven that the set of maps implemented by neural networks with a ReLU activation function is identical to the set of piecewise linear continuous maps. Furthermore, such networks induce a hyperplane arrangement splitting the input domain of the network into convex polyhedra $G_{J}$ over which a network $\Phi$ operates in an affine manner.
In this work, we leverage these properties to define an equivalence class $\sim_{\Phi}$ on top of an input dataset, which can be split into two sets related to the local rank of $\Phi_{J}$ and the intersections $\cap\text{Im}\Phi_{J_{i}}$. We refer to the latter as the overlap decomposition $\mathcal{O}_{\Phi}$ and prove that if the intersections between each polyhedron and an input manifold are convex, the homology groups of neural representations are isomorphic to quotient homology groups $H_{k}(\Phi(\mathcal{M}))\simeq H_{k}(\mathcal{M}/\mathcal{O}_{\Phi})$. This lets us intrinsically calculate the Betti numbers of neural representations without the choice of an external metric. We develop methods to numerically compute the overlap decomposition through linear programming and a union-find algorithm.
Using this framework, we perform several experiments on toy datasets showing that, compared to standard persistent homology, our overlap homology-based computation of Betti numbers tracks purely topological rather than geometric features. Finally, we study the evolution of the overlap decomposition during training on several classification problems while varying network width and depth and discuss some shortcomings of our method.
@@ 1 Introduction
Deep learning is an incredibly powerful framework for learning and generalizing to highly complex tasks. Despite its widespread success, our theoretical understanding of this framework is still lacking. What we do know is that deep neural networks are universal approximators (Hornik, (1991)) and therefore have unbounded expressive power. If we further restrict ourselves to neural networks with ReLU activation functions, we also know that they are exactly equal to the set of continuous piecewise-linear (CPWL) functions (Arora et al., (2018)).
@@
A CPWL function applies a different affine function over different pieces of its domain. Naturally, this phenomenon is also observed in neural networks, where the structure of their weights splits the domain into different polyhedra called linear regions (Pascanu et al., (2013); Montufar et al., (2014)). Quantifying the number (Arora et al., (2018); Raghu et al., (2017); Serra et al., (2018); Hanin and Rolnick, (2019)) as well as other properties (Fan et al., (2023)) of the linear regions that a network generates is thought to reflect the expressivity of a neural network and is a rich subfield in its own right.
@@
Another avenue of research that has shown promise in many fields of science is Topological Data Analysis (TDA), which attempts to quantify the topological features of data (Carlsson, (2009); Wasserman, (2018)). In the context of deep learning, it has been used to relate the topology of training trajectories (Birdal et al., (2021); Dupuis et al., (2023)) and network weights (Rieck et al., (2018); Gabrielsson and Carlsson, (2019); Gutiérrez-Fandiño et al., (2021); Andreeva et al., (2023)) to the generalization error of a neural network. In addition, while counting linear regions is one measure of expressivity, another way to quantify it is to look at the topology of the decision boundary of a neural network (Guss and Salakhutdinov, (2018); Petri and Leitão, (2020); Grigsby and Lindsey, (2022)) or the degree to which a network can change the topology of the input domain as it is propagated through its layers (Naitzat et al., (2020); Wheeler et al., (2021)).
@@
TDA includes a plethora of methods, among which are persistence landscapes (Bubenik et al., (2015)), persistence images (Adams et al., (2017)), persistent laplacians (Mémoli et al., (2022)) and the Mapper algorithm (Singh et al., (2007)). These and most methods in TDA are based on the concept of persistence and more specifically - the persistent homology method (Zomorodian and Carlsson, (2004)). Despite the power of this method, it is known that it is sensitive to outliers, undersampling, and highly non-linear mappings. Furthermore, it is based on calculating distances between points and therefore requires the choice of a metric. This choice leads to the identification of not only topological features, but also geometric ones like curvature (Bubenik et al., (2020)) and convexity (Turkes et al., (2022)). That might be a benefit in some cases, but when reasoning solely about topological features, we would like to avoid contamination from geometric sources.
@@
These issues are especially relevant in the context of deep neural networks as they apply highly non-linear transformations of the input domain that increase curvature (Poole et al., (2016)) in the output space, making their analysis incredibly difficult. To avoid this, we study topological features in the input space through a purely topological lens. We do this using arguments from quotient and relative homology (Hatcher, (2002)), which except for a few works (Pokorny et al., (2016); Blaser and Brun, (2022); Beshkov and Einevoll, (2024)), have received surprisingly little attention in the deep learning literature.
@@ 1.1 Our contributions
Instead of dealing with an external space induced by a map, in which the choice of a metric is ambiguous, we can imagine the following. Take a dataset, determine which pieces are glued to each other by the map, and then count the holes in the manifold after this gluing. Under some simple assumptions, these three steps can be rigorously identified with a manifold $\mathcal{M}$, an equivalence class $\sim_{\Phi}$ and the quotient homology groups $H_{k}(\mathcal{M}/\sim_{\Phi})$ respectively (see Appendix A for background on these terms). This procedure avoids the leakage of geometric information into our estimates of homology since it only depends on the quotient space $\sim_{\Phi}$. It does require knowledge or at least a good estimate of the topology of the manifold $\mathcal{M}$ and a reliable way to identify $\sim_{\Phi}$.
In this work, we show how these steps can be carried out in the context of ReLU neural networks, further developing previous work relating polyhedral decompositions and homology (Liu et al., (2023)). There are two difficulties with making such an approach rigorous. Firstly, how can we actually calculate the equivalence class $\sim_{\Phi}$? Secondly, given this class, how and when can we calculate the quotient homology groups $H_{k}(\mathcal{M}/\sim_{\Phi})$? Our fundamental realization is that since ReLU neural networks split the input domain into convex polyhedra, we can separate the sources of gluing (non-injectivity) into the local rank of the map at each polyhedron and the non-trivial intersections between the images of different polyhedra. We refer to these as the rank and the overlap source, respectively, and show that the latter can be computed exactly, using feasibility checks solved through linear programming. Afterwards, we prove under which conditions the overlap decomposition is sufficient to calculate quotient homology groups.
We compare the performance of persistent and quotient homology when it comes to determining the Betti numbers on toy datasets with a known topology. Our approach gives a new perspective on how topological information propagates through the layers of a neural network. Using quotient homology, we observe that topological simplification happens much more gradually than observed by previous persistent homology-based calculations (Naitzat et al., (2020)). Finally, we study the properties of the overlap decomposition in randomly initialized and trained neural networks and show that the volume of overlapping regions decreases after training, whereas the number of overlap regions tends to increase.
@@ 2 Decompositions of Neural Networks
[…]
@@ 2.1 ReLU networks and linear regions
As shown in Arora et al., (2018), every ReLU neural network is equivalent to a continuous piecewise-linear function and vice versa. Such a network decomposes input space into convex polyhedra (Montufar et al., (2014); Balestriero et al., (2019); Zhang and Wu, (2020); Liu et al., (2023)). This happens as a result of the fact that the affine functions $T_{l}$ can be interpreted as specifying conditions $CT_{l}$, where $C$ is some diagonal matrix with entries in $\{-1,1\}$ defining a polyhedron $P=\{x|CT_{l}(x)\leq 0\}$. Each subsequent layer further decomposes the input space into more polyhedra, since it adds more conditions to the previous decomposition.
@@
Then the local layer decomposition supported on $R$ is defined by the set $\mathcal{L}^{l}|_{R}=\{L^{l}_{J}|_{R}|\forall J\}$.
While the codewords within the same region of the local layer decomposition are constant, that does not mean that the network applies the same affine function over the region (see B.6 in the Appendix). However, this does hold in the polyhedral decomposition (Liu et al., (2023)). Instead of looking at the codeword at layer $l$ we can stack all codeword vectors from the previous layers $C_{l}(x)=[c_{1}(x),...,c_{l}(x)]$ forming a new global codeword vector. This leads us to the next decomposition,
@@
Then the polyhedral decomposition supported on $R$ is defined by the set going over all possible $J$’s or $\mathcal{G}^{l}|_{R}=\{G^{l}_{J}|_{R}|\forall J\}$.
@@ 2.2 The overlap decomposition
The polyhedral decomposition is known in the literature, here we define a new type of decomposition that can be leveraged to compute the homology groups of a neural network, which we will call the overlap decomposition. Our approach is intuitively justified by the realization that if neural networks are equivalent to continuous piecewise-linear functions, then the only way that the topology of the input space can change is through non-injective transformations. Since the network operates piecewise, we can have non-injectivity from two sources; see Appendix B.2 for a proof: $H_{k}([-\pi,\pi]^{2}/\mathcal{O}_{\Phi})$
@@
    The map $\Phi^{l}_{J}$ is low-rank and projects the region $G^{l}_{J}$ to a lower-dimensional subspace. We call this the rank source.
@@
    For a set of maps $\{\Phi^{l}_{J_{1}},...,\Phi^{l}_{J_{N}}\}$ on different polyhedra, there is a non-trivial intersection $\bigcap\limits_{n}\text{Im}\Phi^{l}_{J_{n}}\neq\emptyset$. We call this the overlap source.
The first condition has been described in the work of Beshkov and Einevoll, (2024), but as we will see later, it can be ignored when the intersections between the data manifold and the regions $G^{l}_{J}$ are convex. Therefore, we will focus on defining the overlap decomposition based on the second condition, with the understanding that one should be careful when this condition fails to hold. With this in mind, let us define the overlap decomposition.
@@
A neural network $\Phi^{l}$ induces an equivalence relation that describes all regions of input space on which it is non-injective. It determines equivalence classes $x]=\{y|\Phi^{l}(x)=\Phi^{l}(y)\}$ which together form the set,
@@
The overlap decomposition describes the set of equivalence classes coming from the intersection source and is given by,
[…]
Note that the equivalence relation in equation 4 is reminiscent of a Reeb space (Edelsbrunner et al., (2008)), with the exception that we do not require that $x$ and $y$ lie on the same connected component $\Phi^{-1}(\Phi(x))=\Phi^{-1}(\Phi(y))$. This definition is most useful for intuition and writing proofs. However, when finding overlaps algorithmically we use a coarser definition, as seen in equation 7, that produces identical results. This coarser decomposition can be fully computed given that we have knowledge of the polyhedral decomposition $\mathcal{G}^{l}$, by identifying the intersections between all pairs of different polyhedra and using a union-find structure on top (Kleinberg, (2006)). In the next section, we discuss how to compute this decomposition for a neural network and a supporting dataset. We also discuss how it can be leveraged to compute the homology groups of a neural representation.
@@ 3 Relating the Overlap Decomposition and Homology
Figure 1: A) Illustration of the steps of our method. The manifold $\mathcal{M}$ is transformed by a network into a space on which the two corners are glued together. This space is homeomorphic to the punctured cylinder below it in which the corners $C_{1}$ and $C_{2}$ have been glued to a point. B) Visualization of our method, the top row shows two example non-linear curves, while the bottom shows their estimated homology groups through persistent homology (green) and quotient homology (black). Points that are glued through the overlap decomposition are given the same color and can be seen in the output space (top row) and the input space (the gray lines in the bottom row). As highlighted by the blue boxes in the second example, our method identifies the points that are glued by the neural network.
3.1 Numerical Determination of the Overlap Decomposition through Linear Programming
We have stated that the overlap decomposition is a subset of the equivalence class $\sim_{\Phi}=\{[x]|\Phi(x)=\Phi(y)\}$. However, we have yet to show how to actually compute this set when working with finite data samples. As we will prove later, under the assumption that the intersection between a polyhedron and the data manifold $P\cap\mathcal{M}$ is a convex set (this is automatically true if $\mathcal{M}$ is convex), we can use the polyhedral decomposition to exactly determine the overlaps between data samples. For two populated polyhedra, meaning there exist points in the dataset $\mathcal{D}$ that live in each of these polyhedra, the overlaps $\Phi_{l}(P_{1})\cap\Phi_{l}(P_{2})$ can be determined through linear programming. This is done by checking if given a point $p\in P_{1}$ there is another point $p^{\prime}\in P_{2}$ such that $\Phi_{l}(p)=\Phi_{l}(p^{\prime})$. This second point $p^{\prime}$ can be found through linear programming under the constraints given by the H-representation of $P_{1}$ and $P_{2}$. Thus, to check whether there is a point in $P_{2}$ that overlaps with $p\in P_{1}$, we get the following feasibility linear program,